a 4m fence casts a shadow that is 3m long. how far is it from the tip of the shadow to the top of the fence. show solution.
Answer by Jk22(389)

the fence is 4m high, hence vertically,
the shadow is 3m long horizontally.
this builds a triangle with a right angle, and cathetes 4 and 3 m resp.
By the pythagorean theorem, the hypothenuse is then : sqrt(3*3+4*4)=sqrt(9+16)=sqrt(25)=5 m